OpenStudy (anonymous):

343 /64 x (4/7)^2 - 5/12

OpenStudy (lena772):

Do you know BODMAS?

OpenStudy (anonymous):

\[\frac{ 343 }{ 64 } \times \left( \frac{ 4 }{ 7 } \right)^{2} - \frac{ 5 }{ 12 }\]

OpenStudy (anonymous):

what does BODMAS stand for @Lena772

OpenStudy (lena772):

Brackets over Division Multiplication addition subtraction

OpenStudy (lena772):

So you solve the equation in that order

OpenStudy (anonymous):

i did i got \[\frac{ 5483 }{ 3124 }\]

OpenStudy (anonymous):

but it is wrong :(

OpenStudy (anonymous):

the answer is somehow 4/5

OpenStudy (lena772):

Did you find the value of the second term before doing anything else?

OpenStudy (anonymous):

you mean the number with the exponent? did i do that first? yes

OpenStudy (anonymous):

@Lena772

OpenStudy (lena772):

Yes thats what i meant. Hmmm

OpenStudy (lena772):

(343/64) * 0.32653061224 - (5/12)

OpenStudy (lena772):

(5.359375*0.32653061224)-(5/12)

OpenStudy (anonymous):

What im lost!!!!!

OpenStudy (lena772):

(1.74999999997)-(5/12)

OpenStudy (lena772):

(1.74999999997)-(0.41666666666)

OpenStudy (anonymous):

where did you get those numbers from?

OpenStudy (the_fizicx99):

GO Lena! ~

OpenStudy (lena772):

All I did was convert them to decimals

OpenStudy (lena772):

Your answer is 1.33333333331. I don't know who told you it was 4/5.

OpenStudy (anonymous):

thats the answer is the book is 4/5

OpenStudy (lena772):

Then put the answer in the book.

OpenStudy (anonymous):

@mathmale can you help?

OpenStudy (lena772):

@tHe_FiZiCx99 can probably help you.

OpenStudy (anonymous):

@Lena772 I hope @tHe_FiZiCx99 can help me.... maybe they know why the answer is 4/5

OpenStudy (the_fizicx99):

\(\ \Large \sf \frac{ 343 }{ 64 } \times \left( \frac{ 4 }{ 7 } \right)^{2} - \frac{ 5 }{ 12 } \) \(\ \sf \dfrac{343}{64} * \dfrac{16}{49} - \dfrac{5}{12} \implies \dfrac{5488}{3136} - \dfrac{5}{12} = \dfrac{65856}{37632} - \dfrac{15680}{37632} = \dfrac{50176}{37632} \) That's a big number u_u" using a calculator you still get 1.333.. 4/5 isn't correct.

OpenStudy (the_fizicx99):

However, this does simplify to 4/3.

OpenStudy (mathstudent55):

\(\dfrac{343}{64} \times \left( \dfrac{4}{7} \right)^{2} - \dfrac{5}{12} \) \(= \dfrac{343}{64} \times \dfrac{16}{49} - \dfrac{5}{12} \) \(= \dfrac{343}{\cancel{64}~4} \times \dfrac{\cancel{16}~1}{49} - \dfrac{5}{12} \) \(= \dfrac{343 \times 1}{4 \times 49} - \dfrac{5}{12} \) \(= \dfrac{343}{196} - \dfrac{5}{12} \) \(= \dfrac{343 \times 3}{196 \times 3} - \dfrac{5 \times 49}{12 \times 49} \) \(= \dfrac{1029}{588} - \dfrac{245}{588} \) \(= \dfrac{784}{588}\) \(= \dfrac{196 \times 4}{196 \times 3}\) \(= \dfrac{4}{3}\)